# The Endless Tuning An Artificial Intelligence Design To Avoid Human Replacement and Trace Back Responsibilities

### v1

@@ Abstract
The Endless Tuning is a design method for a reliable deployment of artificial intelligence based on a double mirroring process, which pursues both the goals of avoiding human replacement and filling the so-called responsibility gap [1]. Originally depicted in [2] and ensuing the relational approach urged therein, it was then actualized in a protocol, implemented in three prototypical applications regarding decision-making processes (respectively: loan granting, pneumonia diagnosis, and art style recognition) and tested with such as many domain experts. Step by step illustrating the protocol, giving insights concretely showing “a different voice”[3] in the ethics of artificial intelligence, a philosophical account of technical choices (e.g., a reversed and hermeneutic deployment of XAI algorithms) will be provided in the present study together with the results of the experiments, focusing on user experience rather than statistical accuracy. Even thoroughly employing deep learning models, full control was perceived by the interviewees in the decision-making setting, while it appeared that a bridge can be built between accountability and liability in case of damage.
00footnotetext: This paper introduces the “Endless Tuning” framework, developed by the author. A code demonstration, including subrepositories for each case study, is available at: https://github.com/eliogrande/TheEndlessTuning.git. Interactive versions can also be executed via Google Colab: https://colab.research.google.com/drive/1m1mDWTlE5egzT_oSR18hHLlcwwrX401N?authuser=1#scrollTo=3pMgbEeXeB5X. This material is released under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License (CC BY-NC-SA 4.0). Commercial applications (including integration into proprietary systems or consulting services) require prior written consent from the author.
@@ 1 Introduction
Much work has already been carried out regarding over-reliant behaviour: ex multis, not to mention the foundational [4] on System 1 and System 2, the persistent character and the sluggish one of our mind, [5] experimentally recorded commission and omission errors due to cognitive laziness in the presence of automated decision aids; a pilot study in [6] observed that explanations in artificial intelligence can well increase reliance on recommendations even when they are incorrect; about to predict whether a person’s annual income would exceed \$50K based on some demographic and job information, [7] noticed that showing confidence scores above 80% significantly enhanced trust through an increase in switch percentage (changing idea); [8] successfully tested three cognitive forcing functions (on-demand explanations, updating decisions, waiting time to obtain explanations) in a decision-making process; in a cooperative spirit, machine learning techniques like learning to reject [9] and to defer [10] suggested a socratic artificial intelligence [11].
@@
While maintaining a general definition of trust as «the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability»[12], what the aforementioned anecdote (which, be it true or not, it suffices to reckon it as plausible) brings to light, even independently of the faithfulness of outcomes, lies instead in the ontological properties attributed to the trustee. Indeed, what is it that really seems to matter far beyond and before the automation bias itself? In a deeper sense, the fact that we bestow on artificial intelligence the ability to tell something, adorning with a «style: impressionism», a «trustworthy applicant: yes», or even a «exchange value: \$15k»what is simply a feature, often not transparently computed. But if an algorithm can say something, it can affirm or deny; if it can affirm or deny, it can judge; if it can judge, it can think and has willingness. A bit too much, isn’t it?
@@
Strictly coupled with the mandate of action given to an algorithm, the fateful question arises: who is going to pay in case of damage — obviously given that a causal connection (or at least a reasonable probabilistic one) between the damage itself and the artificial intelligence system is proved? Such a noble concept as the European HLEG’s requirement of human agency and oversight [13], be it meant as human-in-the-loop (see f.e. [14]) or, in reverse, a machine-in-the-loop paradigm (see, under certain respects, [15]), could be not so effective until the so-called responsibility gap [1] is filled, i.e. until the artificial agent’s agency gets hooked to human responsibility. Yet, how to achieve this goal? Not surprisingly, brotherhood comes about among the artificial intelligence ethicist, the computer scientist (or the developer) and the jurist (or the lawyer) — the first one becoming the link between the other two figures.
@@
Starting by focusing on «what acting means», [2] underlines the absence of awareness and willingness in artificial agency, reporting the distortion produced by some anthropomorphic projection. Nonetheless, stumbling just a step outside the perimeter of this projection, breaching the theoretical sphere and literally staying out of sight, machine learning algorithms act. Despite, for example, remaining blind and deterministic from a mathematical viewpoint, the neurons of a neural network shooting layer by layer cease to appear so blind with respect to the meaning or, better, the sense of the activations themselves. Again in the act of activating. A veracious and tangible way to ethics of artificial intelligence could precisely capitalize on this discrepancy, in the effort to get in touch, to establish a relation [16, 17] between different modes of acting. We put forward, indeed, that a design that deals with this relation, traversed back and forth and covering the various entities involved (the producer, the developer, the interface designer, the end-user, and so on), might well constitute the intersection point between accountability and liability.
@@
A comprehensive study concerning the criteria of imputability with respect to artificial intelligence and civil liability [18] shows that currently, in the European Union’s law, a complementary approach is looked for between the ex ante and ex post dimensions of responsibility, where «to the first one corresponds a goal of ascription of responsibility concerning the subjects that produce and use AI systems in the context of their activities»— the EU Artificial Intelligence Act [19], specifying requirements and obligations for providers and operators according to the risk implied toward fundamental human rights — «while the second deals with completing the protections’ framework set up by the safety legislation, providing protection to damaged people also with respect to risks not considered by this last one»[18, p. 321] [^1] Translation is mine. — the new Product Liability Directive for defective products [20], with some space left for national law concerning objective liability. Due to the complexity of the supply chain (known as the “problem of many hands” [21]) and the partially unpredictable conduct of the algorithms themselves, the rules of accountability can serve the purpose of better distributing responsibility [18, pp. 322 ff.], whereas liability usually focuses more on the entity of the damage to be repaired.
@@
In medias res, partly entering and partly building the previously mentioned relation, we propose the Endless Tuning, an ethical design method of artificial intelligence that, at least with regard to decision-making processes (classification or regression tasks), allows to employ artificial intelligence models of nearly any depth avoiding the replacement of the user — but correspondingly requiring them of being competent — and still having the possibility of tracing back responsibilities. The general method itself, which has already been depicted in [2], will be given further insights in Section 2. Secondly, a protocol has been built upon this method, which will be presented in Section 3, step by step explaining its modules and the reasons thereof. In the third place, the protocol itself was implemented in three prototypical versions (that is, so as many apps), according to three different tasks (loan granting, art style recognition, and chest x-ray pneumonia diagnosis), and it was tested through interviewing three domain experts (see A,B and C for transcriptions). Section 4 will provide technical details of the implemented prototypes and sum up the results of the interviews. Section 5 will further discuss the calibration of trust and mandate of action and reflect on whether and how the responsibility gap gets filled.
@@
Practical responses should be given to practical issues; but, even within the framework of fundamental rights, not by imposing a specific ethical viewpoint embedded in a model or a system: rather than ethics, that would be moralism. Similarly to a pro-ethical design [22, p. 190], the Endless Tuning does instead provide the operator with the freedom to fairly deploy an artificial intelligence system. Such freedom is not meant to be merely bounded by the properties of the system itself (positively, thanks to the GUI’s layout, rather than negatively, by the means of data-driven biases [23]), rather resembling a rail on which actions are forced to run. Concerning the term “interface”, [24, p. 7] underlines how the suffix “inter-” before the substantive “face” indicates properties of commonality, link and relation among human beings and between and with real or virtual artifacts. «It is necessary to reflect on the fact that the interface does not exist and it is just an abstract entity which we refer to […]» [^2] Translation is mine. Delving into such an idea, an interface would appear to us as the way in which a relation reveals itself.
@@
Figure 1: A graphical representation of the double loop of Endless Tuning (image taken from [2]; originally made by the same author of this paper; received permission by the editors).
@@
A sort of Continual Learning [25] is in fact depicted here, composed of a double negative feedback. Since the present article concentrates on ethical design, concerning proper machine learning challenges and possible objections thereabout not only might it be worth noting that Continual Learning is a growing field, but also that here we are proposing machine learning on single (or few) tasks, carrying out finetuning by the means of simple rehearsal techniques, which can prove themselves effective. Yet, without referring to specific architectures, we would like to go straight to one point. Our issue certainly lies in contrasting each other’s biases, but when a human learner/decision-maker comes up beside a machine learning model, such artificial intelligence device emerges as an action’s center of gravity. Such a human learner, getting involved and having to balance choices in the meanwhile of interacting, gives the learning process a real time shade, also dealing with an even subtle behavioral deviation margin. The idea to reciprocally profit from each other’s fragility was conceptually illustrated in [2] — after all, the moral, cognitive, physical and even metaphysical fragility of human beings has been notorious since the dawn of time, while for example the media sensation about the COMPAS software [26, 27] publicly disclosed an Achilles heel of artificial intelligence. Sections 3 and Section 4 will dive into the actual implementation. Yet, a ground point remains to be clarified: is there really a different voice to be heard? [3] If machine learning algorithms reflect us as mirrors [28, 29], can they mirror themselves in us? Or is instead the Endless Tuning itself based on a false friend, victim of a silly deceit [30]?
@@
Despite being of utmost significance, accuracy itself might not be the only key. As a statistical measure, accuracy is strictly dependent on the size and the internal variety of the learning data sample — a shift in the data distribution might be lethal in inference mode and, as for the training phase, produce a sudden loss in performance called catastrophic forgetting (see, f. e., [31, 25]). Additionally, we usually employ artificial intelligence one inference at a time. Moreover, outside the laboratory setting, it is reasonable to find out that people use artificial intelligence precisely when they ignore the ground truth. Still, except for transparent models like decision trees and so on, it remains difficult to clearly explain even a single inference because «internal states of machine learning models are non-concepts, concepts that have no intuitive equivalent in the real world and that can only be represented in terms of what they are not»[32].
@@
Such “vulnerability” could instead become a pivotal resource. Little by little, passing through a spectrum of presence not meant as subjectivity, an artificial intelligence algorithm might disclose itself, if not as an interpreter, at least as a useful prompter. Take a vanilla neural network as a guinea pig, not even a deep one, simply trained in supervised fashion. Does it sensibly emerge as an entropy reducer, even if such entropy is not computable? Or, if you prefer, does it return nonsensical truths as ancient sophists distressed Plato [33] by telling sensible lies?
@@
Suppose the model predicts something. Strictly speaking, a feature vector enters the network making neurons consequently “spike” in a merely deterministic fashion. Our lack of knowledge about the internal process could nonetheless lead to unforeseen and sometimes enigmatic results. Tricky questions: how artificial is such an incomprehensible artifact? So, why are we ignorant about that? It is not solely a matter of complexity. Let us take a step back. Our model was trained according to a meta-algorithm (backpropagation) and through a meaningful (for us) target. It is us, for example, who connect the neuron of the last layer indexed 2 with "dog". Consequently, while its obscurity conceptually increases, it comes out that its function has some meaning beyond a strictly numerical one. On the one hand, information gets inscribed in a digital object. Simondon would perhaps have called it a transductive process, carried out «by a mechanism that is analogous to that of perception in the living», to build up a transducer [34, pp. 155-156]. On the other hand, not only is this transductive process (loss and gradient calculation, optimizer steps) itself unreachable even recording and reviewing all the steps with naked eye, but no tool exists to catch how the backpropagation or the inference treats that other sphere of sense beyond a numerical one. Right within this crevice (which could come out to be a rift), neither able to know, nor to think, nor to tell anything, artificial intelligence acts, somehow beyond categories: its acting is without action [35, pp. 38-40].
@@
On a similar path, Romele [36, pp. 89 ff.] suggests the concepts of “imaginative machines” and “emagination” — a form of productive imagination which, sometimes obscurely, embeds the schematizations of our mind in a digital device. [^3] At the time of [2], I had not had the occasion of reading this book. I would like to mention it now. To my pleasure and surprise, a double loop similar to the one of the Endless Tuning was depicted there (p. 104). Despite some differences, a somehow common spirit drove us to a similar intuition. Drawing from sociology and hermeneutics, it talks about «a new form of complex mediation, no longer reducible to a single cognitive system»[36, p. 98]: the new algorithmic machines, nowadays all imaginative at a certain level of abstraction, taking a break from human imagination. [36] mentions [37, p. 2]: algorithmic opacity above all «stems from the mismatch between mathematical optimization in high-dimensionality characteristic of machine learning and the demands of human-scale reasoning and styles of semantic interpretation». It might not be news to say that our vanilla neural network has or even is a self-standing structure, but it is not at all obvious to suggest it is a sort of reactive one. Our goal is to land, say, on the dark side of such opacity: let us move a last step before passing to the protocol.
@@
A crystalline germ containing «no more than an extremely small amount of energy» gets in touch with an amorphous body endowed with potential energy: «its structure and its orientation take control of this energy of the metastable state» and the crystallization starts [38, pp. 80 ff.] Mutatis mutandis, such crystalline germs remind us of the phenomenon of adversarial examples, which in [2] were philosophically suspected to be bugs in the nature of artificial intelligence itself, since an usable object seems to become unusable precisely because of its usability. Adversarial examples are data instances affected by extremely small perturbations ($x-x^{\prime}<r$), which are capable of profoundly (and, paradoxically, precisely) altering the outcome or the explanation (see e.g. [39]) of a model with respect to the original data. For example, a picture that with naked eye is perfectly recognizable as one of a school bus may well be classified with high confidence as an ostrich. Starting from [40], discovering them and calling them «somewhat universal», much work has been carried out trying to understand their nature, typology, construction methods and defenses (ex multis, [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]).
@@
Far from seeking solutions to such mathematically complex problems, we are merely gathering insights. While [50] or [54] focus on a curse of data dimensionality, reporting the ease of moving a point outside the bounds designed by the classifier in a high-dimensionality setting, [53] showed the existence of useful but non-robust features — a standard model trained on a non-robust, mislabeled dataset was nonetheless capable of generalizing on a standard test set. Being useful but imperceptible, such features appeared peculiarly meaningful for the model. Yet, [53] attributed the “adversarial” issue to a possible misalignment between the inherent geometry of data and the l-norm metric employed to build the adversarial examples. Disagreeing with a strongly geometrical approach, [55] underlines that adversarial examples could be affected by features that are «neither signal nor random noise», calling them “artifacts”, or «systematic, reproducible patterns in transformed signals that are created by interactions between our instruments and the world», sort of “entities of mind” just like Doppler effects. Be it as it may, such errors exist, say, oscillating between being a formal and a material issue. As reading Romele would suggest, Levinas [56] would perhaps have noticed:
@@
Perhaps we have found some levers to lift a relation, say, beyond accuracy and inaccuracy and somehow outside taxonomies of autonomous [57] and moral machines [58]. As Section 2 has been focusing on “the trace of the other” in machine learning, we refer to Section 3 concerning the need for a reverse employment of eXplainable Artificial Intelligence algorithms in a hermeneutic fashion. Despite being very far from admitting personality, dignity or consequently rights in computers, the approach of Endless Tuning calls for the government rather than the strict control of artificial intelligence systems, asymmetrically recognizing reciprocal fragility. If we were to define such a relation, we would borrow from Levinas the expression “a contact without contact”. All in all, it is a soft tie, despite the amount of computational resources needed and the risks involved. Paradoxically, “dissecting the (nonexistent) soul” of machines seems the only way to avoid ending up on the couch of an artificial psychoanalyst.

### v3

@@ Abstract
The Endless Tuning is a design method for a reliable deployment of artificial intelligence based on a double mirroring process, which pursues both the goals of avoiding human replacement and filling the so-called responsibility gap [matthias2004responsibility]. Originally depicted in [fabris2024towards] and ensuing the relational approach urged therein, it was then actualized in a protocol, implemented in three prototypical applications regarding decision-making processes (respectively: loan granting, pneumonia diagnosis, and art style recognition) and tested with such as many domain experts. Step by step illustrating the protocol, giving insights concretely showing “a different voice”[gilligan1993different] in the ethics of artificial intelligence, a philosophical account of technical choices (e.g., a reversed and hermeneutic deployment of XAI algorithms) will be provided in the present study together with the results of the experiments, focusing on user experience rather than statistical accuracy. Even thoroughly employing deep learning models, full control was perceived by the interviewees in the decision-making setting, while it appeared that a bridge can be built between accountability and liability in case of damage.
00footnotetext: This paper introduces the “Endless Tuning” framework, developed by the author. A code demonstration, including subrepositories for each case study, is available at: https://github.com/eliogrande/TheEndlessTuning.git. Interactive versions can also be executed via Google Colab: https://colab.research.google.com/drive/1m1mDWTlE5egzT_oSR18hHLlcwwrX401N?authuser=1#scrollTo=3pMgbEeXeB5X. This material is released under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License ($C$ BY-NC-SA 4.0). Commercial applications (including integration into proprietary systems or consulting services) require prior written consent from the author.
@@ 1 Introduction
Much work has already been carried out regarding over-reliant behaviour: ex multis, not to mention the foundational [kahneman2011thinking] on System 1 and System 2, the persistent character and the sluggish one of our mind, [skitka1999does] experimentally recorded commission and omission errors due to cognitive laziness in the presence of automated decision aids; a pilot study in [bansal2021does] observed that explanations in artificial intelligence can well increase reliance on recommendations even when they are incorrect; about to predict whether a person’s annual income would exceed \$50K based on some demographic and job information, [zhang2020effect] noticed that showing confidence scores above 80% significantly enhanced trust through an increase in switch percentage (changing idea); [buccinca2021trust] successfully tested three cognitive forcing functions (on-demand explanations, updating decisions, waiting time to obtain explanations) in a decision-making process; in a cooperative spirit, machine learning techniques like learning to reject [cortes2016learning] and to defer [madras2018predict] suggested a socratic artificial intelligence [punzi2023towards].
@@
While maintaining a general definition of trust as «the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability»[lee2004trust], what the aforementioned anecdote (which, be it true or not, it suffices to reckon it as plausible) brings to light, even independently of the faithfulness of outcomes, lies instead in the ontological properties attributed to the trustee. Indeed, what is it that really seems to matter far beyond and before the automation bias itself? In a deeper sense, the fact that we bestow on artificial intelligence the ability to tell something, adorning with a «style: impressionism», a «trustworthy applicant: yes», or even a «exchange value: \$15k»what is simply a feature, often not transparently computed. But if an algorithm can say something, it can affirm or deny; if it can affirm or deny, it can judge; if it can judge, it can think and has willingness. A bit too much, isn’t it?
@@
Strictly coupled with the mandate of action given to an algorithm, the fateful question arises: who is going to pay in case of damage — obviously given that a causal connection (or at least a reasonable probabilistic one) between the damage itself and the artificial intelligence system is proved? Such a noble concept as the European HLEG’s requirement of human agency and oversight [hleg2019ethics], be it meant as human-in-the-loop (see f.e. [wu2022survey]) or, in reverse, a machine-in-the-loop paradigm (see, under certain respects, [miller2023explainable]), could be not so effective until the so-called responsibility gap [matthias2004responsibility] is filled, i.e. until the artificial agent’s agency gets hooked to human responsibility. Yet, how to achieve this goal? Not surprisingly, brotherhood comes about among the artificial intelligence ethicist, the computer scientist (or the developer) and the jurist (or the lawyer) — the first one becoming the link between the other two figures.
@@
Starting by focusing on «what acting means», [fabris2024towards] underlines the absence of awareness and willingness in artificial agency, reporting the distortion produced by some anthropomorphic projection. Nonetheless, stumbling just a step outside the perimeter of this projection, breaching the theoretical sphere and literally staying out of sight, machine learning algorithms act. Despite, for example, remaining blind and deterministic from a mathematical viewpoint, the neurons of a neural network shooting layer by layer cease to appear so blind with respect to the meaning or, better, the sense of the activations themselves. Again in the act of activating. A veracious and tangible way to ethics of artificial intelligence could precisely capitalize on this discrepancy, in the effort to get in touch, to establish a relation [Buber1970-BUBIAT, fabris2016relazione] between different modes of acting. We put forward, indeed, that a design that deals with this relation, traversed back and forth and covering the various entities involved (the producer, the developer, the interface designer, the end-user, and so on), might well constitute the intersection point between accountability and liability.
@@
A comprehensive study concerning the criteria of imputability with respect to artificial intelligence and civil liability [dal2024intelligenza] shows that currently, in the European Union’s law, a complementary approach is looked for between the ex ante and ex post dimensions of responsibility, where «to the first one corresponds a goal of ascription of responsibility concerning the subjects that produce and use AI systems in the context of their activities»— the EU Artificial Intelligence Act [EU2024AIAct], specifying requirements and obligations for providers and operators according to the risk implied toward fundamental human rights — «while the second deals with completing the protections’ framework set up by the safety legislation, providing protection to damaged people also with respect to risks not considered by this last one»[dal2024intelligenza, p. 321] [^1] Translation is mine. — the new Product Liability Directive for defective products [EU2024PLD], with some space left for national law concerning objective liability. Due to the complexity of the supply chain (known as the “problem of many hands” [nissenbaum1996accountability]) and the partially unpredictable conduct of the algorithms themselves, the rules of accountability can serve the purpose of better distributing responsibility [dal2024intelligenza, pp. 322 ff.], whereas liability usually focuses more on the entity of the damage to be repaired.
@@
In medias res, partly entering and partly building the previously mentioned relation, we propose the Endless Tuning, an ethical design method of artificial intelligence that, at least with regard to decision-making processes (classification or regression tasks), allows to employ artificial intelligence models of nearly any depth avoiding the replacement of the user — but correspondingly requiring them of being competent — and still having the possibility of tracing back responsibilities. The general method itself, which has already been depicted in [fabris2024towards], will be given further insights in Section 2. Secondly, a protocol has been built upon this method, which will be presented in Section 3, step by step explaining its modules and the reasons thereof. In the third place, the protocol itself was implemented in three prototypical versions (that is, so as many apps), according to three different tasks (loan granting, art style recognition, and chest x-ray pneumonia diagnosis), and it was tested through interviewing three domain experts (see A,B and C for transcriptions). Section 4 will provide technical details of the implemented prototypes and sum up the results of the interviews. Section 5 will further discuss the calibration of trust and mandate of action and reflect on whether and how the responsibility gap gets filled.
@@
Practical responses should be given to practical issues; but, even within the framework of fundamental rights, not by imposing a specific ethical viewpoint embedded in a model or a system: rather than ethics, that would be moralism. Similarly to a pro-ethical design [floridi2014fourth, p. 190], the Endless Tuning does instead provide the operator with the freedom to fairly deploy an artificial intelligence system. Such freedom is not meant to be merely bounded by the properties of the system itself (positively, thanks to the GUI’s layout, rather than negatively, by the means of data-driven biases [ntoutsi2020bias]), rather resembling a rail on which actions are forced to run. Concerning the term “interface”, [zannoni2024design, p. 7] underlines how the suffix “inter-” before the substantive “face” indicates properties of commonality, link and relation among human beings and between and with real or virtual artifacts. «It is necessary to reflect on the fact that the interface does not exist and it is just an abstract entity which we refer to […]» [^2] Translation is mine. Delving into such an idea, an interface would appear to us as the way in which a relation reveals itself.
@@
Figure 1: A graphical representation of the double loop of Endless Tuning (image taken from [fabris2024towards]; originally made by the same author of this paper; received permission by the editors).
@@
A sort of Continual Learning [lesort2020continual] is in fact depicted here, composed of a double negative feedback. Since the present article concentrates on ethical design, concerning proper machine learning challenges and possible objections thereabout not only might it be worth noting that Continual Learning is a growing field, but also that here we are proposing machine learning on single (or few) tasks, carrying out finetuning by the means of simple rehearsal techniques, which can prove themselves effective. Yet, without referring to specific architectures, we would like to go straight to one point. Our issue certainly lies in contrasting each other’s biases, but when a human learner/decision-maker comes up beside a machine learning model, such artificial intelligence device emerges as an action’s center of gravity. Such a human learner, getting involved and having to balance choices in the meanwhile of interacting, gives the learning process a real time shade, also dealing with an even subtle behavioral deviation margin. The idea to reciprocally profit from each other’s fragility was conceptually illustrated in [fabris2024towards] — after all, the moral, cognitive, physical and even metaphysical fragility of human beings has been notorious since the dawn of time, while for example the media sensation about the COMPAS software [larson2016compas, angwin2016machinebias] publicly disclosed an Achilles heel of artificial intelligence. Sections 3 and Section 4 will dive into the actual implementation. Yet, a ground point remains to be clarified: is there really a different voice to be heard? [gilligan1993different] If machine learning algorithms reflect us as mirrors [fabris2012etica, valera2022espejos], can they mirror themselves in us? Or is instead the Endless Tuning itself based on a false friend, victim of a silly deceit [natale2022macchine]?
@@
Despite being of utmost significance, accuracy itself might not be the only key. As a statistical measure, accuracy is strictly dependent on the size and the internal variety of the learning data sample — a shift in the data distribution might be lethal in inference mode and, as for the training phase, produce a sudden loss in performance called catastrophic forgetting (see, f. e., [chen2018continual, lesort2020continual]). Additionally, we usually employ artificial intelligence one inference at a time. Moreover, outside the laboratory setting, it is reasonable to find out that people use artificial intelligence precisely when they ignore the ground truth. Still, except for transparent models like decision trees and so on, it remains difficult to clearly explain even a single inference because «internal states of machine learning models are non-concepts, concepts that have no intuitive equivalent in the real world and that can only be represented in terms of what they are not»[offert2017know].
@@
Such “vulnerability” could instead become a pivotal resource. Little by little, passing through a spectrum of presence not meant as subjectivity, an artificial intelligence algorithm might disclose itself, if not as an interpreter, at least as a useful prompter. Take a vanilla neural network as a guinea pig, not even a deep one, simply trained in supervised fashion. Does it sensibly emerge as an entropy reducer, even if such entropy is not computable? Or, if you prefer, does it return nonsensical truths as ancient sophists distressed Plato [plato_sophist_1986] by telling sensible lies?
@@
Suppose the model predicts something. Strictly speaking, a feature vector enters the network making neurons consequently “spike” in a merely deterministic fashion. Our lack of knowledge about the internal process could nonetheless lead to unforeseen and sometimes enigmatic results. Tricky questions: how artificial is such an incomprehensible artifact? So, why are we ignorant about that? It is not solely a matter of complexity. Let us take a step back. Our model was trained according to a meta-algorithm (backpropagation) and through a meaningful (for us) target. It is us, for example, who connect the neuron of the last layer indexed 2 with "dog". Consequently, while its obscurity conceptually increases, it comes out that its function has some meaning beyond a strictly numerical one. On the one hand, information gets inscribed in a digital object. Simondon would perhaps have called it a transductive process, carried out «by a mechanism that is analogous to that of perception in the living», to build up a transducer [alma99244708110502171, pp. 155-156]. On the other hand, not only is this transductive process (loss and gradient calculation, optimizer steps) itself unreachable even recording and reviewing all the steps with naked eye, but no tool exists to catch how the backpropagation or the inference treats that other sphere of sense beyond a numerical one. Right within this crevice (which could come out to be a rift), neither able to know, nor to think, nor to tell anything, artificial intelligence acts, somehow beyond categories: its acting is without action [fabris2002paradossi, pp. 38-40].
@@
On a similar path, Romele [romele2019digital, pp. 89 ff.] suggests the concepts of “imaginative machines” and “emagination” — a form of productive imagination which, sometimes obscurely, embeds the schematizations of our mind in a digital device. [^3] At the time of [fabris2024towards], I had not had the occasion of reading this book. I would like to mention it now. To my pleasure and surprise, a double loop similar to the one of the Endless Tuning was depicted there (p. 104). Despite some differences, a somehow common spirit drove us to a similar intuition. Drawing from sociology and hermeneutics, it talks about «a new form of complex mediation, no longer reducible to a single cognitive system»[romele2019digital, p. 98]: the new algorithmic machines, nowadays all imaginative at a certain level of abstraction, taking a break from human imagination. [romele2019digital] mentions [burrell2016machine, p. 2]: algorithmic opacity above all «stems from the mismatch between mathematical optimization in high-dimensionality characteristic of machine learning and the demands of human-scale reasoning and styles of semantic interpretation». It might not be news to say that our vanilla neural network has or even is a self-standing structure, but it is not at all obvious to suggest it is a sort of reactive one. Our goal is to land, say, on the dark side of such opacity: let us move a last step before passing to the protocol.
@@
A crystalline germ containing «no more than an extremely small amount of energy» gets in touch with an amorphous body endowed with potential energy: «its structure and its orientation take control of this energy of the metastable state» and the crystallization starts [gilbert2020individuation, pp. 80 ff.] Mutatis mutandis, such crystalline germs remind us of the phenomenon of adversarial examples, which in [fabris2024towards] were philosophically suspected to be bugs in the nature of artificial intelligence itself, since an usable object seems to become unusable precisely because of its usability. Adversarial examples are data instances affected by extremely small perturbations ($x-x^{\prime}<r$), which are capable of profoundly (and, paradoxically, precisely) altering the outcome or the explanation (see e.g. [razmi2023interpretation]) of a model with respect to the original data. For example, a picture that with naked eye is perfectly recognizable as one of a school bus may well be classified with high confidence as an ostrich. Starting from [szegedy2013intriguing], discovering them and calling them «somewhat universal», much work has been carried out trying to understand their nature, typology, construction methods and defenses (ex multis, [goodfellow2014explaining, yuan2019adversarial, zhang2019adversarial, zhao2017generating, xie2019information, li2021verifying, madry2017towards, moosavi2017universal, kurakin2018adversarial, chattopadhyay2019curse, elsayed2018adversarial, hendrycks2021natural, ilyas2019adversarial, shafahi2018adversarial, buckner2020understanding]).
@@
Far from seeking solutions to such mathematically complex problems, we are merely gathering insights. While [chattopadhyay2019curse] or [shafahi2018adversarial] focus on a curse of data dimensionality, reporting the ease of moving a point outside the bounds designed by the classifier in a high-dimensionality setting, [ilyas2019adversarial] showed the existence of useful but non-robust features — a standard model trained on a non-robust, mislabeled dataset was nonetheless capable of generalizing on a standard test set. Being useful but imperceptible, such features appeared peculiarly meaningful for the model. Yet, [ilyas2019adversarial] attributed the “adversarial” issue to a possible misalignment between the inherent geometry of data and the l-norm metric employed to build the adversarial examples. Disagreeing with a strongly geometrical approach, [buckner2020understanding] underlines that adversarial examples could be affected by features that are «neither signal nor random noise», calling them “artifacts”, or «systematic, reproducible patterns in transformed signals that are created by interactions between our instruments and the world», sort of “entities of mind” just like Doppler effects. Be it as it may, such errors exist, say, oscillating between being a formal and a material issue. As reading Romele would suggest, Levinas [levinas1963trace] would perhaps have noticed:
@@
Perhaps we have found some levers to lift a relation, say, beyond accuracy and inaccuracy and somehow outside taxonomies of autonomous [tosic2004towards] and moral machines [moor2006nature]. As Section 2 has been focusing on “the trace of the other” in machine learning, we refer to Section 3 concerning the need for a reverse employment of eXplainable Artificial Intelligence algorithms in a hermeneutic fashion. Despite being very far from admitting personality, dignity or consequently rights in computers, the approach of Endless Tuning calls for the government rather than the strict control of artificial intelligence systems, asymmetrically recognizing reciprocal fragility. If we were to define such a relation, we would borrow from Levinas the expression “a contact without contact”. All in all, it is a soft tie, despite the amount of computational resources needed and the risks involved. Paradoxically, “dissecting the (nonexistent) soul” of machines seems the only way to avoid ending up on the couch of an artificial psychoanalyst.